# Beyond Prompts: Dynamic Conversational Benchmarking of Large Language Models

### v1

@@ 1 Introduction
In response, we introduce the LTM Benchmark [^1] All data, experiments, and code for implementing, and running the LTM benchmark mentioned in this paper are under an open source license and available, in full, at: https://github.com/GoodAI/goodai-ltm-benchmark. This project is licensed under the MIT License., an automated system designed to evaluate the Long-Term Memory (LTM) and Continual Learning (CL) capabilities of conversational agents. The LTM Benchmark engages agents in a single, prolonged conversation, incorporating multiple tasks and distractions to simulate realistic and meaningful interactions. This approach provides a more comprehensive assessment of an agent’s ability to effectively use and integrate information across extended dialogues. Furthermore, we show that this “conversational multitasking” structure significantly degrades the test performance of LLMs, indicating that their real-world capabilities are not fully revealed through most contemporary benchmarks. The LTM benchmark is primarily synthetic, and can be generated to contain conversations of arbitrary length, which may potentially surpass the context size of the LLMs under test.
@@ 3 The LTM benchmark
As a mostly synthetic benchmark, the generation processes is both transparent and reproducible. As such, we have released all benchmark results and generators. The benchmark is dynamic, but all generation and scheduling processes are deterministic. The one source of nondeterminism in the testing is from the agents themselves. The length of agents’ replies will have knock-on effects in the amount of filler tokens that are used, and that in turn may influence the replies of the agent. More details on this and our mitigation strategies can be found in the supplementary material.
@@ 3.1 Test structure
The conversation begins with a message explaining the overall situation to the agent, and continues with all tests’ messages and agent responses (see example in Figure 3). Every test encompasses a series of interventions from the benchmark system (taking the user’s role), which are uniformly distributed across a previously-delimited area in the conversation known as the memory span of the test (see Figure 2). The test’s span can be set independently from the agent’s effective input size, which allows for tests surpassing current LLMs’ context lengths. Other constraints like token and time waits (see episodic memory and Jokes scenario in §3.2) also form part of the test information, which is provided to the benchmark’s scheduling system.
@@
A test can be arbitrarily complex depending on its content and definition. Test’s messages are either delivered in a static manner (providing a script) or dynamically (via a functional description), depending on the requirements of the testing scenario. The LTM benchmark’s scheduling system coordinates the execution of tests and weaves them together into a seamless conversation, while respecting the memory span, test compatibility constraints and optimizing the use of tokens. Finally, once the last test question is answered, the test is scored according to the agent’s performance. The complexity of this scoring mechanism can vary greatly too, from a simple pattern matching statement to an elaborate evaluation protocol using LLMs as evaluators. The supplementary material contains more details relating to the content and evaluations methods of the tests.
@@ 4 Results
We have produced a standard configuration of the LTM benchmark and use it to evaluate and compare some of the current leading LLMs in addition to a baseline LTM system. For the configuration, we have included three repetitions of each scenario, and we have set the number of needles for the scenarios requiring it: in Colours, three changes; in Name List, five different names; in Jokes, four jokes; in Locations Directions, the path goes through six locations; and in Shopping List the list updates six times. We run the same benchmark with a number of different memory spans: 2k, 32k, 120k, 200k, and 500k tokens. Additionally, we include results for the same tasks evaluated in isolated scenarios, where there is neither task-switching nor distractors in the conversation (resembling standard benchmarks). We also exclude ChapterBreak from the 2k configuration to avoid exceeding the memory span.
Table 1: Benchmark results for distinct agent configurations: open-source LLMs, commercial LLMs, and LTM agents. Context stands for the maximum input tokens available to the LLM. All tests are out of 11 points, apart from the 2k benchmark, which is out of 10. A visual representation of these results can be seen in Figure 5.
@@
Mixtral (8x7B) 32000 5.0 2.0 1.4 0.6 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1
Mixtral (8x22B) 65536 4.9 2.5 5.6 1.6 0.0 0.0 0.1 0.1 0.1 0.1 0.1 0.1
Llama 3 (70B) 8000 8.2 1.1 1.9 2.4 0.2 0.0 0.2 0.0 0.2 0.0 0.2 0.0
GPT-3.5-Turbo 16384 4.1 1.6 4.7 1.6 0.1 0.1 0.0 0.0 0.0 0.0 0.0 0.0
GPT-4-Turbo 128000 7.9 0.1 6.6 1.0 4.8 3.9 5.8 2.4 3.9 1.8 1.0 0.8
GPT-4o 128000 7.6 2.0 5.9 1.2 4.6 3.6 5.5 2.6 5.2 1.7 0.9 0.6
Claude-Opus 200000 8.3 1.4 7.8 1.8 6.7 2.3 7.4 2.7 5.4 1.7 3.4 0.6
Gemini 1.5 Pro 1000000 7.4 2.2 6.5 1.5 6.4 1.7 7.0 0.9 8.0 2.8 5.3 2.6
LTM (Opus) 16384 8.7 1.9 7.5 1.4 5.0 2.3 5.7 3.4 6.4 2.9 4.9 2.5
LTM (GPT-4) 16384 9.2 0.7 6.3 1.7 5.2 1.9 5.0 2.4 5.3 1.9 3.1 1.6
LTM (Llama3 (70B)) 8000 8.4 1.6 6.9 2.3 5.0 2.2 4.7 2.2 5.6 1.3 4.8 1.5
Each benchmark run consists of 33 tests (11 scenarios and 3 repetitions). We average the results corresponding to repetitions of the same test type, and compute a standard deviation out of those scores. In total, an agent can achieve up to one point per scenario, which makes for a maximum score of 11 points.
[…]
Our baseline LTM agent [^2] Details of the LTM system can be found on GitHub at: https://github.com/GoodAI/goodai-ltm. uses a vector database for storing memories, which it queries before every response. Additionally, it has a JSON scratchpad which it can use for keeping volatile or dynamically-changing information, and utilizes query rewriting for the vector database. The agent uses an LLM as a core controller, retrieving up to 10 relevant memories on each turn, and using up to 16k tokens of the LLM’s context [^3] We only use fewer than 16k tokens with Llama 3, which has an input limit of 8k tokens..
In Table 1 we show the scores achieved by each agent configuration under all memory span settings. We also show a visual representation of the same results in Figure 5. These results show that LLMs perform well on short memory spans, but they struggle as soon as the memory span surpasses their maximum context size, as the relevant information is simply lost. Conversely, the LTM agent suffers less when the memory span is increased.
Figure 5: Scores obtained for all agent and memory span configurations. Solid bars correspond to vanilla LLMs, while strided bars represent LLMs used as a core controller for the LTM agent setup. Each color can be associated with a different LLM. Context sizes for each of the agents vary, and are detailed in Table 1.
@@ 5 Analysis
Looking at the results in detail (see the supplementary material), the Restaurant task was the most variable one. Claude had trouble with the task by being too expressive, adding “stage directions” to its responses. The Mixtral models also often hallucinated interactions between the waiter and diner on their own, resulting in failures. A lot of models that otherwise fail completely on the other benchmark tasks, can get some credit for the first part of the Restaurant task where they are greeted, shown a menu and asked for a drink order in a single message. If the agent plays along with the scenario, it will obtain some score. Our automated evaluation system (powered in some parts by GPT-4 turbo) also fails at some points, like being unable to distinguish between dishes that are discussed, and ones that have been ordered.
@@ 7 Limitations and future work
Agents’ robustness in this paper are based on only three repetitions, which is not very reliable. The benchmark has a high time and financial cost to running it, and therefore multiple repeats have proven currently infeasible. Also, our episodic memory testing sometimes uses time jumps, which require the agent to run on the same machine as the benchmark system.
Given the breadth of possible responses to any question, the automatic evaluation systems can perform poorly in some edge cases, requiring manual checking. Summarized reports facilitate the manual revision, but in the future we aim to make evaluations infallible. See the supplementary material for examples where the automated checking has failed, and how we have manually evaluated the tests.
Future work includes addressing the noted shortcomings and adding more tests, which include but are not limited to 1.) Learning and recalling instructions or facts, especially focusing on forward and backward transfer; 2.) Stress testing with large amounts of information to integrate, 3.) More research-based tasks where multiple passages need to be recalled and integrated to answer correctly; and 4.) Multi-modal evaluations.
@@ 8 Conclusions
In this paper, we identified an issue with the current suite of benchmarks for LLMs, namely that the tests they run do not necessarily reflect real usage as chat agents along with the complexities that a chat environment brings. In response to this we have presented a new benchmark that subjects conversational agents (LLM-based or not) to a series of tests over a single lengthy conversation in order to test their memory and ability to integrate information. We benchmark current SOTA large-context LLMs, as well as LLMs equipped with a basic LTM system.

### v2

@@ 1 Introduction
In response, we introduce the LTM Benchmark [^1] All data, experiments, and code for implementing, and running the LTM benchmark mentioned in this paper are under an open source license and available, in full, at: https://github.com/GoodAI/goodai-ltm-benchmark., an automated system designed to evaluate the Long-Term Memory (LTM) and Continual Learning (CL) capabilities of conversational agents. The LTM Benchmark engages agents in a single, prolonged conversation, incorporating multiple tasks and distractions to simulate realistic and meaningful interactions. This approach provides a more comprehensive assessment of an agent’s ability to effectively use and integrate information across extended dialogues. Furthermore, we show that this “conversational multitasking” structure significantly degrades the test performance of LLMs, indicating that their real-world capabilities are not fully revealed through most contemporary benchmarks. The LTM benchmark is primarily synthetic, and can be generated to result in conversations of arbitrary length, which may potentially surpass the context size of the LLMs under test.
@@ 3 The LTM benchmark
As a mostly synthetic benchmark, the generation processes is both transparent and reproducible. As such, we have released all benchmark results and generators. The benchmark is dynamic, but all generation and scheduling processes are deterministic. The one source of nondeterminism in the testing is from the agents themselves. The length of agents’ replies will have knock-on effects in the amount of filler tokens that are used, and that in turn may influence the replies of the agent. More details on this and our mitigation strategies can be found in Appendix C.
@@ 3.1 Test structure
The conversation begins with a message explaining the overall situation to the agent, and continues with all tests’ messages and agent responses (see Figure 3). Every test encompasses a series of interventions from the benchmark system (taking the user’s role), which are uniformly distributed across a previously-delimited area in the conversation known as the memory span of the test (see Figure 2). The test’s span can be set independently from the agent’s effective input size, which allows for tests surpassing current LLMs’ context lengths. Other constraints like token and time waits (see episodic memory and Jokes scenario in §3.2) also form part of the test information, which is provided to the benchmark’s scheduling system.
@@
A test can be arbitrarily complex depending on its content and definition. Test’s messages are either delivered in a static manner (providing a script) or dynamically (via a functional description), depending on the requirements of the testing scenario. The LTM benchmark’s scheduling system coordinates the execution of tests and weaves them together into a seamless conversation, while respecting the memory span, test compatibility constraints and optimizing the use of tokens. Finally, once the last test question is answered, the test is scored according to the agent’s performance. The complexity of this scoring mechanism can vary greatly too, from a simple pattern matching statement to an elaborate evaluation protocol using LLMs as evaluators. Appendix A contains more details relating to the content and evaluations methods of the tests.
@@ 3.2 Test scenarios
3.3 Generation processes
This benchmark is considered mainly synthetic because of how the data is presented, where tasks are interwoven with other tasks and previous agent responses, and not so much because the data itself being synthetic. Some of the test scenarios do not rely on synthetic data, but on already existing –and sometimes relatively small– datasets. It is the interleaving of tasks and messages what makes each interaction unique, since the conversation history is highly diverse and sensitive to the agent’s responses.
Each test scenario has its own generation method, which is conditioned on the global seed as well as the index of the test repetition. All data required for running the tests is generated upfront during initialization and saved to disk. Dynamic tests also rely on seeded random generators and zero-temperature LLM calls to synthesize data and react to the agent’s responses. We refer the reader to Appendix A for a detailed description of the generation of each scenario included in this edition of the LTM benchmark.
Besides this initial data generation process, the benchmark relies on a scheduling system to coordinate the intertwined delivery of test messages while attending to the different tests’ requirements and other constraints. In order to accomplish that, each test is required to communicate those constraints along with every message delivered: namely, how much time it should pass or how many tokens should be exchanged before its next intervention, if any at all. This is similar to how yielding mechanisms work in computational multithreading. The policy of this message scheduling system can be summarized by the following rules:
  • Incompatible tests cannot run concurrently.
  • Messages from a same test are delivered until the test ends or yields.
  • When a test yields or ends, the system gives priority to resuming other waiting tests before starting a new one.
  • Only if no test can be resumed or started, the system will try to unblock the situation by sending dummy tasks (using tokens) or performing time jumps (using up time).
A detailed specification of the benchmark’s scheduling system and policy can be found in Appendix C.
3.4 LTM Implementations
In these tests, we test both unmodified LLMs, as well as LLMs augmented with LTM systems. These LTM implementations fall into broadly two categories: those that implement memory through modifications of the LLM, and those that implement memory through modifications to the context of the LLM, but not through architectural choices.
Works such as LongMem [Wang et al., 2023] capture the hidden states of the transformer and use those residuals as input to a “SideNet” which can attend to a cached memory of inputs or documents. Memorizing Transformers [Wu et al., 2022] augment one of the final layers of a transformer model with a $k$-nn search over hidden states saved to a stack. MemoryLLM [Wang et al., 2024] is another model which maintains a pool of updatable parameters within the latent space of the transformer.
Works such as MemoryBank [Zhong et al., 2024] and MemGPT [Packer et al., 2023] instead develop memory as a tool which augments the input context of the LLM, rather than influencing token generation directly. These approaches allow for some flexibility in their choice of model, and permit the use of commercial LLMs such as Claude or GPT.
Agents with LTMs can be further distinguished by whether they use their LTM actively or passively. In an active system, the LLM has some awareness of the functionality of the LTM, and can choose how to use it. This often takes the form of a multi-step resolution to a user query by making decisions such as when memory should be read from or written to. Passive systems on the other hand perform reads and writes automatically in a typically fixed loop.
With the aim of offering some reference results for agents with LTM, in our tests we include results from our baseline LTM system, as well as from MemoryBank and MemGPT. Other models were not included due to a lack of pretrained weights or a reliance on early completion LLMs, which are not instruction fine-tuned and cannot withstand the required conversational interaction.
MemoryBank places the focus on empathy and psychological companionship, and stores conversation and summaries of those conversations in chronological order, along with a summary of the user. This information is kept in a vector database, and is passively queried during its conversation with the user. MemoryBank also implements a human-like memory forgetting curve.
An example of an active system is MemGPT. This agent attempts to answer queries by consulting multiple forms of memory. It contains an in-context core memory, and external archival and recall memories. The core memories are always visible as a part of the context, whereas the archival and recall memories, need to be queried and are for storing long running events or documents, and the chat history respectively. The agent can update its archival or core memories via function calls.
Our baseline LTM system [^2] More details of the LTM system can be found on GitHub at: https://github.com/GoodAI/goodai-ltm. uses both a vector database and a JSON scratchpad. On receiving a user message, the agent generates multiple queries for the vector database and retrieves chunks that are not currently part of the context. The scratchpad is updated (or not) after the LLM generates a response to the query, and is similar to MemGPT’s core memory in that it is always present in the context.
@@ 4 Results
We have produced a standard configuration of the LTM benchmark and used it to evaluate and compare some of the current leading LLMs, in addition to a series of LTM-enabled agents. For the configuration, we have included three repetitions of each scenario, and we have set the number of needles for the scenarios requiring it: in Colours, three changes; in Name List, five different names; in Jokes, four jokes; in Locations Directions, the path goes through six locations; and in Shopping List the list updates six times. We run the same benchmark with a number of different memory spans: 2k, 32k, 120k, 200k, and 500k tokens. Additionally, we include results for the same tasks evaluated in isolated scenarios, where there is neither task-switching nor distractors in the conversation (resembling standard benchmarks). We also exclude ChapterBreak from the 2k configuration to avoid exceeding the memory span.
Table 1: Benchmark results for distinct agent configurations: open-source LLMs, commercial LLMs, and LTM agents. Context stands for the maximum input tokens available to the LLM. All tests are out of 11 points, apart from the 2k benchmark, which is out of 10. Underscored values are the result of two benchmark runs. A visual representation of these results can be seen in Figure 5.
@@
Mixtral 8x7B 32000 5.0 0.9 1.4 0.5 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1
Mixtral 8x22B 65536 4.9 1.0 5.6 0.7 0.0 0.0 0.1 0.1 0.1 0.1 0.1 0.1
Llama 3 70B 8000 8.2 0.7 1.9 1.0 0.2 0.0 0.2 0.0 0.2 0.0 0.2 0.0
GPT-3.5 turbo 16384 4.1 0.8 4.7 0.8 0.1 0.1 0.0 0.0 0.0 0.0 0.0 0.0
GPT-4 turbo 128000 7.9 0.1 6.6 0.7 5.2 1.3 5.8 1.0 3.9 0.9 1.0 0.5
GPT-4o 128000 7.6 0.9 5.9 0.7 5.6 1.4 5.9 1.1 5.2 0.9 0.9 0.5
GPT-4o-mini 128000 7.8 0.7 5.6 1.0 4.1 0.7 5.1 1.2 4.8 0.8 1.4 0.6
Claude 3 Opus 200000 8.3 0.8 7.8 0.9 6.7 1.0 7.4 1.1 5.1 0.9 3.4 0.4
Gemini 1.5 Pro 1000000 7.4 0.9 6.5 0.8 6.4 0.8 7.0 0.5 7.7 1.0 6.1 1.3
LTM Claude 3 Opus 16384 8.7 0.9 7.5 0.7 5.0 1.0 5.7 1.2 6.4 1.1 4.9 1.0
LTM GPT-4 turbo 16384 9.2 0.5 6.3 0.8 5.2 0.9 5.0 1.0 5.3 0.9 3.1 0.8
LTM GPT-4o-mini 16384 8.1 0.7 5.4 0.8 4.7 1.0 4.6 1.1 4.2 0.9 4.7 1.0
LTM Llama 3 70B 8000 8.4 0.8 6.9 1.0 5.0 0.9 4.7 1.0 5.6 0.7 4.8 0.8
MemGPT GPT-4o-mini 16384 - 40000 6.8 0.9 4.8 0.7 2.3 0.7 3.2 0.8 3.6 0.8 2.0 0.7
MemoryBank GPT-4o-mini 16384 - 60000 6.4 0.7 3.5 0.7 2.0 0.5 2.8 0.5 3.3 0.8 3.6 0.6
A benchmark run consists of 33 tests (11 scenarios and 3 repetitions each), which measure the agent’s performance on the different scenarios under diverse contexts. All scores are normalized to $[0,1]$ and scores from the same scenario are averaged together, which makes for a maximum of 11 points per benchmark. In order to obtain accurate distribution estimates, we generate 1,000 possible outcomes by repeatedly sampling one result from each scenario and adding them together. After running the initial set of benchmarks for all configurations, we allocated the remaining budget to configurations where reducing variance would be most impactful —such as those with high scores and competing closely with others. These additional runs use a different seed and are marked with underscores in Table 1.
The benchmarks are executed using a variety of on-demand APIs from various providers, including Google, OpenAI, and Anthropic. The open source models were tested via TogetherAI’s API. Running and evaluating all the benchmarks in this paper cost $\approx$ \$7,200 and took 142 hours.
The tested LTM agents are capable of using different LLMs. On a cost-to-quality ratio, we found that GPT-4o-mini works well. We have tried to keep the context limit for the LTM agents to 16k tokens [^3] We only use fewer than 16k tokens with Llama 3, which has an input limit of 8k tokens., but the MemoryBank and MemGPT systems cannot artificially restrict their context sizes. We observed that the volume of memories which they retrieve later in the benchmark can balloon the context sizes to around 60k or 40k tokens respectively.
In Table 1 we show the scores achieved by each agent configuration under all memory span settings. We also show a visual representation of the same results in Figure 5. These results show that LLMs perform well on short memory spans, but they struggle as soon as the memory span surpasses their maximum context size, as the relevant information is simply lost. Conversely, LTM agents suffer less when the memory span is increased.
Figure 5: Scores obtained for all agent and memory span configurations. Solid boxes correspond to vanilla LLMs, while strided boxes represent LTM agents. Each color can be associated with a different LLM. Context sizes for each of the agents vary, and are detailed in Table 1.
@@ 5 Analysis
Looking at the results in detail (see Appendix D), the Restaurant task was the most variable one. Claude had trouble with the task by being too expressive, adding “stage directions” to its responses. The Mixtral models also often hallucinated interactions between the waiter and diner on their own, resulting in failures. A lot of models that otherwise fail completely on the other benchmark tasks, can get some credit for the first part of the Restaurant task where they are greeted, shown a menu and asked for a drink order in a single message. If the agent plays along with the scenario, it will obtain some score. Our automated evaluation system (powered in some parts by GPT-4 turbo) also fails at some points, like being unable to distinguish between dishes that are discussed, and ones that have been ordered.
@@ 7 Limitations and future work
Agents’ robustness in this paper are based on only three repetitions, which is not very reliable. The benchmark has a high time and financial cost to running it, and therefore multiple repeats have proven currently infeasible, but we expect this to change in the near future.
Given the breadth of possible responses to any question, the automatic evaluation systems can perform poorly in some edge cases, requiring manual checking. Summarized reports facilitate the manual revision, but in the future we aim to make evaluations infallible. See Appendix B for examples where the automated checking has failed, and how we have manually evaluated the tests.
Future work includes addressing the noted shortcomings and adding more tests, which include but are not limited to 1.) Learning and recalling instructions or facts, especially focusing on forward and backward transfer; 2.) Stress testing with large amounts of information to integrate, 3.) More research-based tasks where multiple passages need to be recalled and integrated to answer correctly; 4.) Multi-modal evaluations; and 5.) Multi-user scenarios, where the agent needs to attend different users’ needs while keeping their information separated.
@@ 8 Conclusions
In this paper, we identified an issue with the current suite of benchmarks for LLMs, namely that the tests they run do not necessarily reflect real usage as chat agents along with the complexities that a chat environment brings. In response to this we have presented a new benchmark that subjects conversational agents (LLM-based or not) to a series of tests over a single lengthy conversation in order to test their memory and ability to integrate information. We benchmark current SOTA large-context LLMs, as well as LLMs equipped with a LTM systems.